# arXiv:2602.22038v2[math.AP] 27 Feb 2026

### v2

@@ 2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}.$ (38)
@@ 2.3 Estimates for $I_{t}$: the nonlinear terms
We now address the nonlinear terms in (21). The main novelty is to use Donsker-Varadhan inequality (82), in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
@@
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
@@
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (41)
We observe that by Donsker-Varadhan inequality (82), for $\eta>>1$
@@
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$
which implies by Csiszár-Kullback-Pinsker inequality (81) and (41)
@@
by Donsker-Varadhan inequality (82), we obtain
@@
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta},$ (45)
for $\eta>>1$.
@@
Now from Leibniz’s rule we deduce, for $\eta>>1$
@@
$\overset{}{\leq}2\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}+2\int_{\mathbb{R}^{d}}\rho_{s}^{N}\left|\nabla\ln\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx,$ (50)
where in the last inequality, we are using Donsker-Varadhan inequality (82):
@@
$+\eta\ln{\left(\int_{\mathbb{R}^{d}}2\exp{(-2\pi|x-X_{t}|^{2})}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx\right)}$
$\overset{}{\leq}\eta\mathcal{H}(\rho_{s}^{N}|\rho_{s})+C_{\eta}.$
[…]
$\left|\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\right|$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle$
$\times\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\right)^{\frac{1}{2}}$
@@
$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\eta}$
$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\eta}.$ (51)
@@
$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\eta}.$ (52)
@@ 2.4 Application of Gronwall’s Lemma
In this subsection, we conclude our estimates, by applying Gronwall’s inequality.
Indeed, with the same computations as in [39], for all $\delta\in(0,\theta_{1})$ there exists a random variable $A_{0}$, finite almost surely, such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{1}+\delta},$
with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
@@
$+C_{\alpha,T}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}$
$+A_{0}N^{-\theta_{1}+\delta}$
$+TC_{\eta}$ (53)
@@
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{2}}+A_{0}N^{-\theta_{1}+\delta}+C,$ (54)
with $\theta_{1}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$ and $\theta_{2}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$.
Finally, from Gronwall’s Lemma in (54), recalling (28), (31) and (32), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+C\right)$ (55)
@@ 2.5 End of the proof: Removing the stopping time
We now remove the stopping time $\tau^{N}$ introduced in (55) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$. These assumptions ensure sufficient control on the initial data and on the particle system, allowing us to pass from the localized estimates to global-in-time bounds with high probability, and finally eliminate the stopping time.

### v3

@@ 2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}N^{-\theta_{1}},$ (46)
for $\theta_{1}\doteq\left(1-\beta\left(1+2/d+2\alpha\right)\right)$.
@@ 2.3 Estimates for $I_{t}$: the nonlinear terms
We now address the nonlinear terms in (29). The main novelty is to use Donsker-Varadhan inequality (108), combined with the notion of solution of limiting process given by Definition 1.1, in the context of moderate interactions, along with the dissipation that comes from of Fisher information and Assumption $(\mathbf{A}^{\nabla\cdot K})$, to tackle with the singularity of kernel $K$.
@@
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\underbrace{\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle}_{\doteq I^{N}}\,dx\,ds$
@@
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|_{\infty}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (49)
We observe that, for $\eta>2C_{1}(8T+C_{3})+1$
$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty$ (50)
which implies by Donsker-Varadhan inequality (108),
@@
Hence by Csiszár-Kullback-Pinsker inequality (107) and (49), we derive
@@
for $\eta>2C_{1,2}(8T+C_{3})+1$, we obtain
$\int_{\mathbb{R}^{d}}\frac{C_{3}}{1+t}\exp{\left(-\frac{|x-X_{t}|^{2}}{8t+C_{3}}\right)}\exp{\left(\frac{C_{1,2}(1+|x-X_{t}|^{2})}{\eta}\right)}\,dx<\infty.$ (55)
It follows by Donsker-Varadhan inequality (108),
@@
Now from Leibniz’s rule we deduce,
@@
$I^{N}$ $\leq 2\|K_{0}\|_{\infty}\left\langle S_{s}^{N},V^{N}(x-\cdot)\right\rangle\left(2\eta\mathcal{H}(\rho^{N}_{s}|\rho_{s})+2\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+C_{\eta}\right)^{\frac{1}{2}}$
@@
$+\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+TC_{\epsilon,\eta}$
$\overset{\text{Holder}}{\leq}(2\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (62)
@@
$I_{t}$ $\leq C_{\epsilon,K_{0},\eta}\int_{0}^{t}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+(3\epsilon+\sqrt{2}/2)\int_{0}^{t}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds+TC_{\epsilon,\eta}.$ (63)
@@ 2.4 Application of Gronwall’s Lemma
In this subsection, we conclude our localizated estimates, by applying Gronwall’s inequality.
Indeed, with the same computations as in [40], for all $\delta\in(0,\theta_{2})$ there exists a random variable $A_{0}$, finite almost surely, such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{2}+\delta},$ (64)
with $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
@@
$+C_{\alpha,T}N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+TC_{\eta}$ (65)
@@
$\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})-\mathcal{H}(\rho_{0}^{N}|\rho_{0})$ $\lesssim\int_{0}^{t\wedge\tau^{N}}\mathcal{H}(\rho_{s}^{N}|\rho_{s})\,ds+N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+T,$ (66)
with $\theta_{1}\doteq 1-\beta(1+\frac{2}{d}+2\alpha)$ and $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
Finally, from Gronwall’s inequality (109) in (66), recalling (36), (39) and (40), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right)$ (67)
@@ 2.5 End of the proof: Removing the stopping time
We now remove the stopping time $\tau^{N}$ introduced in (67) by exploiting the probabilistic framework provided by Assumptions $\left(\mathbf{A}^{\rho_{0}}\right)$ and $\left(\mathbf{A}^{X_{0}}\right)$. It allows us to pass from the localized estimates to uniform bounds with high probability, removing the stopping time.